Find the x and y intercepts for each line. Apply elimination or subsitution to find the coordinates fo the point of the intersection
15x+11y=-17
8x-6y=74

Answer by richwmiller(17219)

15x+11y=-17
8x-6y=74
x=4 y=-7
check
32-+42=74
ok
60-77=-17
ok
90x=66y=-102
88x-66y=814
178x=712
x=4
32-6y=74
-6y=42
y=-7